# Continuous Invariance Learning

### v1

@@ 2.2 The Risks of Existing Methods on Continuous Domain Tasks
Existing approximation methods propose to learn a featurizer by matching $\mathbb{E}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ for different $\boldsymbol{t}$. However, in the continuous environment setting, there are a lot of domains and each domain contains limited amount of samples, leading to noisy empirical estimations $\hat{\mathbb{E}}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$. Specifically, existing methods employ positive losses on each environment $\boldsymbol{t}$ to assess whether $\hat{\mathbb{E}}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ matches the other environments. However, since the estimated $\hat{\mathbb{E}}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ can deviate significantly from $\mathbb{E}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$, these methods fail to identify invariant features. In this part, we first use REx as an example to theoretically illustrate this issue and then show experimental results for other variants.
Theoretical analysis of REx on continuous domain tasks. Consider a simplified case where $\boldsymbol{x}$ is a concatenation of a spurious feature $\boldsymbol{x}_{s}$ and an invariant feature $\boldsymbol{x}_{v}$, i.e., $\boldsymbol{x}=[x_{s},x_{v}]$. Further, $\Phi$ is a feature mask, i.e., $\Phi\in\{0,1\}^{2}$. Let $\Phi_{s}$ and $\Phi_{v}$ denote the feature mask that merely select spurious and invariant feature, i.e., $\Phi_{s}=[1,0]$ and $\Phi_{v}=[0,1]$, respectively. Let ^$\hat{\mathcal{R}}^{t}(\boldsymbol{w},\Phi)=\frac{1}{n^{t}}\sum_{i=1}^{n_{e}}\ell(\boldsymbol{w}(\Phi(\boldsymbol{x}_{i})),y_{i})$ denote the finite sample loss on domain $t$ where $n^{t}$ is the sample size in domain $t$. Denote ^$\hat{\mathcal{L}}_{\text{REx}}(\Phi)$ as the finite-sample REx loss in Eq. (1). We omit the subscript ‘REx’ when it is clear from the context in this subsection. REx can identify the invariant feature mask $\Phi_{v}$ if and only if ^$\hat{\mathcal{L}}(\Phi_{v})<\hat{\mathcal{L}}(\Phi),\forall\Phi\in\{0,1\}^{2}.$ Suppose the dataset $\mathcal{S}$ that contains $|\mathcal{E}|$ domains. For simplicity, we assume that there are equally $n^{t}={n}/{|\mathcal{E}|}$ samples in each domain.
@@
The expected loss of the model using spurious features in each domain, $\mathcal{R}^{t}(\boldsymbol{w},\Phi_{s})$, follows a Gaussian distribution, i.e., $\mathcal{R}^{t}(\boldsymbol{w},\Phi_{s})\sim\mathcal{N}(\mathcal{R}(\boldsymbol{w},\Phi_{s}),\delta_{R})$, where the mean $\mathcal{R}(\boldsymbol{w},\Phi_{s})$ is the average loss over all domains. Given a feature mask $\Phi\in\{\Phi_{s},\Phi_{v}\}$, the loss of an individual sample $(\boldsymbol{x},y)$ from environment $t$ deviates from the expectation loss (of domain $t$) by a Gaussian, $\ell(y,\boldsymbol{w}(\Phi(\boldsymbol{x})))-\mathbb{E}^{t}\left[\ell(y,\boldsymbol{w}(\Phi(\boldsymbol{x})))\right]\sim\mathcal{N}(0,\sigma^{2}_{\Phi})$. The variance $\sigma_{\Phi}$ is drawn from a hyper exponential distribution with density function $p(\sigma_{\Phi};\lambda)=\lambda\exp(-\lambda\sigma_{\Phi})$ for each $\Phi$.
Remark on the Setting and Assumption. Suppose there are infinite samples in each domain, the loss in domain $t$ is $\mathcal{R}^{t}(\boldsymbol{w},\Phi)$. If $\Phi=\Phi_{v}$, $\mathcal{R}^{t}(\boldsymbol{w},\Phi)$ is the same for all domains, leading to zero penalty. If $\Phi=\Phi_{s}$, the loss in domain $t$ deviates from the average loss by a variance according to the assumption, which induces a penalty that scales linearly with the number of environments. Hence, in this case, REx can easily identify the invariant feature mask since the spurious feature mask will have a much larger penalty. This is probably the minimal setting to show whether a method can learn invariance. Our following theoretical results show that, even in such a simple setting, REx can fail if there are many environments and the sample size in each environment is limited. Note that based on this minimal setting, it is easy to construct failure cases for REx for more complex settings.
@@
Proof Sketch and Main Motivation. The complete proof is included in Appendix F. When there are only a few samples in each domain (i.e., $|\mathcal{T}|$ is comparable to $n$), the empirical loss ^$\hat{\mathcal{R}}^{t}(\boldsymbol{w},\Phi)$ deviates from its expectation by a Gaussian variable $\epsilon_{t}~{}\sim~{}\mathcal{N}(0,\sigma_{\Phi}^{2}/n_{t})$. After taking square, we have $\mathbb{E}[\epsilon_{t}^{2}]~{}=~{}\sigma_{\Phi}^{2}/n_{t}$. There are $|\mathcal{T}|$ domains and $n_{t}~{}=~{}n/|\mathcal{T}|$. So we have $\sum_{t}\mathbb{E}[\epsilon_{t}^{2}]=|\mathcal{T}|^{2}\sigma_{\Phi}^{2}/n$, indicating that the data randomness can induce an estimation error that is quadratic in $|\mathcal{T}|$. Note that the expected penalty (assuming we have infinite samples in each domain) grows linearly in $|\mathcal{T}|$. Therefore the estimation error dominates the empirical loss $\hat{\mathcal{L}}$ with large $|\mathcal{T}|$, which means the algorithm selects features based on the data noise rather than the invariance property. Intuitively, existing IRM methods try to find a feature $\Phi$ which aligns $\hat{\mathbb{E}}^{t}[Y|\Phi(\boldsymbol{x})]$ among different $t$. Due to the finite-sample estimation error, $\hat{\mathbb{E}}^{t}[Y|\Phi(\boldsymbol{x})]$ can be far away from $\mathbb{E}^{t}[Y|\Phi(\boldsymbol{x})]$, leading to the failure of invariance learning. One can also show that other variants, e.g., IRMv1, also suffer from this issue (see empirical verification later in this section).
@@
When we have a continuous domain in real-world applications, each sample $(\boldsymbol{x}_{i},y_{i})$ has a distinct domain index, i.e., there is only one sample in each domain. Therefore when $|\mathcal{T}|\approx n$, REx can fail with probability 1/4 when $n\geq\left(\frac{\sigma_{R}}{\delta\mathcal{G}^{-1}(1/4)}\right)^{2}$.
@@ 3 Our Method
Formulation. Suppose $\Phi(\boldsymbol{x})$ successfully extracts invariant features, and we have $\boldsymbol{y}\perp\boldsymbol{t}|\Phi(\boldsymbol{x})$ according to (Arjovsky et al., 2019). The previous analysis shows that it is difficult to align $\mathbb{E}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ for each domain $\boldsymbol{t}$ since there are only very limited samples in $\boldsymbol{t}$ in the continuous domain tasks. This also explains why most existing methods fail in continuous domain tasks. In this part, we propose to align $\mathbb{E}^{y}[\boldsymbol{t}|\Phi(\boldsymbol{x})]$ for each class $\boldsymbol{y}$ (see Appendix D for the discussion on the relationship between $\mathbb{E}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ and $\mathbb{E}^{y}[\boldsymbol{t}|\Phi(\boldsymbol{x})]$). Since there is a sufficient number of samples in each class $\boldsymbol{y}$, we can obtain an accurate estimation of $\mathbb{E}^{y}[\boldsymbol{t}|\Phi(\boldsymbol{x})]$. We perform the following steps to verify whether $\mathbb{E}^{y}[\boldsymbol{t}|\Phi(\boldsymbol{x})]$ is the same for each $\boldsymbol{y}$ : we first fit a function $h\in\mathcal{H}$ to predict $\boldsymbol{t}$ based on $\Phi(\boldsymbol{x})$. Since $\boldsymbol{t}$ is continuous, we adopt L2 loss to measure the distance between $\boldsymbol{t}$ and $h(\Phi(\boldsymbol{x}))$, i.e., $\mathbb{E}[\|h(\Phi(\boldsymbol{x}))-\boldsymbol{t}\|_{2}^{2}]$. We use another function $g$ to predict $\boldsymbol{t}$ based on $\Phi(\boldsymbol{x})$ and $\boldsymbol{y}$, and minimize $\mathbb{E}\|g(\Phi(\boldsymbol{x}),\boldsymbol{y})-\boldsymbol{t}\|_{2}^{2}$. If $\boldsymbol{y}\perp\boldsymbol{t}|\Phi(\boldsymbol{x})$ holds, $\boldsymbol{y}$ does not bring more information to predict $\boldsymbol{t}$ when conditioned on $\Phi(\boldsymbol{x})$. Then the loss achieved by $g(\Phi(\boldsymbol{x}),\boldsymbol{y})$ would be similar to the loss achieved by $h(\Phi(\boldsymbol{x}))$, i.e., $\mathbb{E}[\|g(\Phi(\boldsymbol{x}),\boldsymbol{y})-\boldsymbol{t}\|_{2}^{2}]=\mathbb{E}[\|h(\Phi(\boldsymbol{x}))-\boldsymbol{t}\|_{2}^{2}]$.
@@
$\min_{\Phi,\boldsymbol{w}}~{}\mathbb{E}_{\boldsymbol{x},\boldsymbol{y}}\ell(\boldsymbol{w}(\Phi(\boldsymbol{x})),\boldsymbol{y})\quad\text{s.t.}\min_{h\in\mathcal{H}}\max_{g\in\mathcal{G}}\mathbb{E}[\|h(\Phi(\boldsymbol{x}))-\boldsymbol{t}\|_{2}^{2}-\|g(\Phi(\boldsymbol{x}),\boldsymbol{y})-\boldsymbol{t}\|_{2}^{2}]=0$
In practice, we can replace the hard constraint in (3) with a soft regularization term as follows:
$\min_{\boldsymbol{w},\Phi,h}\max_{g}~{}Q(\boldsymbol{w},\Phi,g,h)=\mathbb{E}_{\boldsymbol{x},\boldsymbol{y},\boldsymbol{t}}\Big{[}\ell(\boldsymbol{w}(\Phi(\boldsymbol{x})),\boldsymbol{y})+\lambda\big{(}\|h(\Phi(\boldsymbol{x})),\boldsymbol{t}\|_{2}^{2}-\|g(\Phi(\boldsymbol{x}),\boldsymbol{y}),\boldsymbol{t}\|_{2}^{2}\big{)}\Big{]},$
@@
Algorithm. We adapt Stochastic Gradient Ascent and Descent (SGDA) to solve (3). SGDA alternates between inner maximization and outer minimization by performing one step of gradient each time. The full algorithm is shown in Appendix A.
@@ 3.1 Theoretical Analysis of Continuous Invariance Learning
In this section, we assume $\boldsymbol{t}$ is a one-dimensional scalar to ease the notation for clarity. All results still hold when $\boldsymbol{t}$ is a vector. We start by stating some assumptions on the capacity of the function classes.
@@
$\mathcal{H}$ contains $h^{*}$, where $h^{*}(\boldsymbol{z}):=\mathbb{E}[\boldsymbol{t}|\boldsymbol{z}=\Phi(\boldsymbol{x})]$.
@@
$\mathcal{G}$ contains $g^{*}$, where $g^{*}(\boldsymbol{z},\boldsymbol{y}):=\mathbb{E}[\boldsymbol{t}|\boldsymbol{z}=\Phi(\boldsymbol{x}),\boldsymbol{y})]$.
@@
Suppose Assumption 2 and 3 hold, $h^{*}$ and $g^{*}$ minimize the following losses given a fixed $\Phi$, we have $h^{*}(\cdot)=\operatorname*{arg\,min}_{h\in\mathcal{H}}\mathbb{E}_{(\boldsymbol{x},t)}(h(\Phi(\boldsymbol{x}))-\boldsymbol{t})^{2},$ and $g^{*}(\cdot)=\operatorname*{arg\,min}_{g\in\mathcal{G}}\mathbb{E}_{(\boldsymbol{x},\boldsymbol{y},t)}(g(\Phi(\boldsymbol{x}),\boldsymbol{y})-\boldsymbol{t})^{2}.$ where the proof is shown in Appendix G.
@@
Suppose Assumptions 2 and 3 hold. The constraint of (3) is satisfied if and only $\mathbb{E}[\boldsymbol{t}|\Phi(\boldsymbol{x})]=\mathbb{E}[\boldsymbol{t}|\Phi(\boldsymbol{x}),\boldsymbol{y}]$ holds for $\forall\boldsymbol{y}\in\mathcal{Y}$.
@@
The advantage of CIL in continuous domain tasks. Consider a 10-class classification task with an infinite number of domains and each domain contains only one sample, i.e., $n\xrightarrow[]{}\infty$, and $n/|\mathcal{T}|=1$, Proposition 1 shows that REx would fail to identify the invariant feature with constant probability. Whereas, Theorem 2 shows that CIL can still effectively extract invariant features. Intuitively, existing methods aim to align $\mathbb{E}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ across different $\boldsymbol{t}$ values. However, in continuous environment settings with limited samples per domain, the empirical estimations $\hat{\mathbb{E}}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ become noisy. These estimations deviate significantly from the true $\mathbb{E}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$, rendering existing methods ineffective in identifying invariant features. In contrast, CIL proposes to align $\mathbb{E}^{y}[\boldsymbol{t}|\Phi(\boldsymbol{x})]$, which can be accurately estimated as there are sufficient samples in each class.
Below we analyze the empirical counterpart of the soft regularization version ((3)) for finite sample performance, i.e.,
$\min_{\boldsymbol{w},\Phi,h}\max_{g}\hat{Q}(\boldsymbol{w},\Phi,g,h):=\hat{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y},\boldsymbol{t}}\Big{[}\ell(\boldsymbol{w}(\Phi(\boldsymbol{x})),\boldsymbol{y})+\lambda\big{(}\|h(\Phi(\boldsymbol{x}))-\boldsymbol{t}\|^{2}_{2}-\|g(\Phi(\boldsymbol{x}),\boldsymbol{y})-\boldsymbol{t}\|_{2}^{2}\big{)}\Big{]}.$
where $\hat{\mathbb{E}}$ is the empirical counterpart of expectation $\mathbb{E}$. Suppose that we solve (3.1) with SGDA (Algorithm 1) and obtain $(\hat{w},\hat{\Phi},\hat{g},\hat{h})$, which is a $(\epsilon_{1},\epsilon_{2})$ optimal solution of (3.1). Specifically, we have
@@
Suppose we solve (3.1) by SGDA as introduced in Algorithm 1 using a training dataset of size $n$ and obtain an $(\epsilon_{1},\epsilon_{2})$ solution $-^(\hat{w},\hat{\Phi},\hat{g},\hat{h})$. Under the assumptions specified in the appendix, we then have with probability at least $1-\delta$ that

### v2

@@ 2.2 The Risks of Existing Methods on Continuous Domain Tasks
Existing approximation methods propose to learn a feature extractor by matching $\mathbb{E}^{t}[\bm{y}|\Phi(\bm{x})]$ for different $\bm{t}$. However, in the continuous environment setting, there are a lot of domains and each domain contains limited amount of samples, leading to noisy empirical estimations $\hat{\mathbb{E}}^{t}[\bm{y}|\Phi(\bm{x})]$. Specifically, existing methods employ positive losses on each environment $\bm{t}$ to assess whether $\hat{\mathbb{E}}^{t}[\bm{y}|\Phi(\bm{x})]$ matches the other environments. However, since the estimated $\hat{\mathbb{E}}^{t}[\bm{y}|\Phi(\bm{x})]$ can deviate significantly from $\mathbb{E}^{t}[\bm{y}|\Phi(\bm{x})]$, these methods fail to identify invariant features. In this part, we first use REx as an example to theoretically illustrate this issue and then show experimental results for other variants.
Theoretical analysis of REx on continuous domain tasks. Consider a simplified case where $\bm{x}$ is a concatenation of a spurious feature $\bm{x}_{s}$ and an invariant feature $\bm{x}_{v}$, i.e., $\bm{x}=[x_{s},x_{v}]$. Further, $\Phi$ is a feature mask, i.e., $\Phi\in\{0,1\}^{2}$. Let $\Phi_{s}$ and $\Phi_{v}$ denote the feature mask that merely select spurious and invariant feature, i.e., $\Phi_{s}=[1,0]$ and $\Phi_{v}=[0,1]$, respectively. Let ^$\hat{\mathcal{R}}^{t}(\bm{w},\Phi)=\frac{1}{n^{t}}\sum_{i=1}^{n_{e}}\ell(\bm{w}(\Phi(\bm{x}_{i})),y_{i})$ denote the finite sample loss on domain $t$ where $n^{t}$ is the sample size in domain $t$. Denote ^$\hat{\mathcal{L}}_{\text{REx}}(\Phi)$ as the finite-sample REx loss in Eq. equation 1. We omit the subscript ‘REx’ when it is clear from the context in this subsection. REx can identify the invariant feature mask $\Phi_{v}$ if and only if ^$\hat{\mathcal{L}}(\Phi_{v})<\hat{\mathcal{L}}(\Phi),\forall\Phi\in\{0,1\}^{2}.$ Suppose the dataset $\mathcal{S}$ that contains $|\mathcal{T}|$ domains. For simplicity, we assume that there are equally $n^{t}={n}/{|\mathcal{T}|}$ samples in each domain.
@@
The expected loss of the model using spurious features in each domain, $\mathcal{R}^{t}(\bm{w},\Phi_{s})$, follows a Gaussian distribution, i.e., $\mathcal{R}^{t}(\bm{w},\Phi_{s})\sim\mathcal{N}(\mathcal{R}(\bm{w},\Phi_{s}),\delta_{R})$, where the mean $\mathcal{R}(\bm{w},\Phi_{s})$ is the average loss over all domains. Given a feature mask $\Phi\in\{\Phi_{s},\Phi_{v}\}$, the loss of an individual sample $(\bm{x},y)$ from environment $t$ deviates from the expectation loss (of domain $t$) by a Gaussian, $\ell(y,\bm{w}(\Phi(\bm{x})))-\mathbb{E}^{t}\left[\ell(y,\bm{w}(\Phi(\bm{x})))\right]\sim\mathcal{N}(0,\sigma^{2}_{\Phi})$. The variance $\sigma_{\Phi}$ is drawn from a hyper exponential distribution with density function $p(\sigma_{\Phi};\lambda)=\lambda\exp(-\lambda\sigma_{\Phi})$ for each $\Phi$.
Remark on the Setting and Assumption. Recall that $\mathcal{R}^{t}(\bm{w},\Phi)$ represents the expected loss of $(\bm{w},\Phi)$ in domain $t$. When $\Phi=\Phi_{v}$, the loss is the same across all domains, resulting in zero penalty. However, when $\Phi=\Phi_{s}$, the loss in domain $t$ deviates from the average loss, introducing a penalty that increases linearly with the number of environments. In this case, REx can easily identify the invariant feature if we have an infinite number of samples in each domain. However, if we have multiple environments with limited sample sizes in each, REx can fail with a constant probability. This limitation will be discussed further in the following parts.
@@
Proof Sketch and Main Motivation. The complete proof is included in Appendix F. When there are only a few samples in each domain (i.e., $|\mathcal{T}|$ is comparable to $n$), the empirical loss ^$\hat{\mathcal{R}}^{t}(\bm{w},\Phi)$ deviates from its expectation by a Gaussian variable $\epsilon_{t}~{}\sim~{}\mathcal{N}(0,\sigma_{\Phi}^{2}/n_{t})$. After taking square, we have $\mathbb{E}[\epsilon_{t}^{2}]~{}=~{}\sigma_{\Phi}^{2}/n_{t}$. There are $|\mathcal{T}|$ domains and $n_{t}~{}=~{}n/|\mathcal{T}|$. So we have $\sum_{t}\mathbb{E}[\epsilon_{t}^{2}]=|\mathcal{T}|^{2}\sigma_{\Phi}^{2}/n$, indicating that the data randomness can induce an estimation error that is quadratic in $|\mathcal{T}|$. Note that the expected penalty (assuming we have infinite samples in each domain) grows linearly in $|\mathcal{T}|$. Therefore the estimation error dominates the empirical loss $\hat{\mathcal{L}}$ with large $|\mathcal{T}|$, which means the algorithm selects features based on the data noise rather than the invariance property. Intuitively, existing IRM methods try to find a feature $\Phi$ which aligns $\hat{\mathbb{E}}^{t}[Y|\Phi(\bm{x})]$ among different $t$. Due to the finite-sample estimation error, $\hat{\mathbb{E}}^{t}[Y|\Phi(\bm{x})]$ can be far away from $\mathbb{E}^{t}[Y|\Phi(\bm{x})]$, leading to the failure of invariance learning. One can also show that other variants, e.g., IRMv1, also suffer from this issue (see empirical verification later in this section).
@@
When we have a continuous domain in real-world applications, each sample $(\bm{x}_{i},y_{i})$ has a distinct domain index, i.e., there is only one sample in each domain. Therefore when $|\mathcal{T}|\approx n$, REx can fail with probability 1/4 when $n\geq\left(\frac{\sigma_{R}}{\delta\mathcal{G}^{-1}(1/4)}\right)^{2}$.
@@ 3 Our Method
Formulation. Suppose $\Phi(\bm{x})$ successfully extracts invariant features, and we have $\bm{y}\perp\bm{t}|\Phi(\bm{x})$ according to Arjovsky et al. (2019). The previous analysis shows that it is difficult to align $\mathbb{E}^{t}[\bm{y}|\Phi(\bm{x})]$ for each domain $\bm{t}$ since there are only very limited samples in $\bm{t}$ in the continuous domain tasks. This also explains why most existing methods fail in continuous domain tasks. In this part, we propose to align $\mathbb{E}^{y}[\bm{t}|\Phi(\bm{x})]$ for each class $\bm{y}$ (see Appendix D for the discussion on the relationship between $\mathbb{E}^{t}[\bm{y}|\Phi(\bm{x})]$ and $\mathbb{E}^{y}[\bm{t}|\Phi(\bm{x})]$). Since there is a sufficient number of samples in each class $\bm{y}$, we can obtain an accurate estimation of $\mathbb{E}^{y}[\bm{t}|\Phi(\bm{x})]$. We perform the following steps to verify whether $\mathbb{E}^{y}[\bm{t}|\Phi(\bm{x})]$ is the same for each $\bm{y}$ : we first fit a function $h\in\mathcal{H}$ to predict $\bm{t}$ based on $\Phi(\bm{x})$. Since $\bm{t}$ is continuous, we adopt L2 loss to measure the distance between $\bm{t}$ and $h(\Phi(\bm{x}))$, i.e., $\mathbb{E}[\|h(\Phi(\bm{x}))-\bm{t}\|_{2}^{2}]$. We use another function $g$ to predict $\bm{t}$ based on $\Phi(\bm{x})$ and $\bm{y}$, and minimize $\mathbb{E}\|g(\Phi(\bm{x}),\bm{y})-\bm{t}\|_{2}^{2}$. If $\bm{y}\perp\bm{t}|\Phi(\bm{x})$ holds, $\bm{y}$ does not bring more information to predict $\bm{t}$ when conditioned on $\Phi(\bm{x})$. Then the loss achieved by $g(\Phi(\bm{x}),\bm{y})$ would be similar to the loss achieved by $h(\Phi(\bm{x}))$, i.e., $\mathbb{E}[\|g(\Phi(\bm{x}),\bm{y})-\bm{t}\|_{2}^{2}]=\mathbb{E}[\|h(\Phi(\bm{x}))-\bm{t}\|_{2}^{2}]$.
@@
$\min_{\Phi,\bm{w}}~{}\mathbb{E}_{\bm{x},\bm{y}}\ell(\bm{w}(\Phi(\bm{x})),\bm{y})\quad\text{s.t.}\min_{h\in\mathcal{H}}\max_{g\in\mathcal{G}}\mathbb{E}[\|h(\Phi(\bm{x}))-\bm{t}\|_{2}^{2}-\|g(\Phi(\bm{x}),\bm{y})-\bm{t}\|_{2}^{2}]=0$
In practice, we can replace the hard constraint in equation 3 with a soft regularization term as follows:
$\min_{\bm{w},\Phi,h}\max_{g}~{}Q(\bm{w},\Phi,g,h)=\mathbb{E}_{\bm{x},\bm{y},\bm{t}}\Big{[}\ell(\bm{w}(\Phi(\bm{x})),\bm{y})+\lambda\big{(}\|h(\Phi(\bm{x}))-\bm{t}\|_{2}^{2}-\|g(\Phi(\bm{x}),\bm{y})-\bm{t}\|_{2}^{2}\big{)}\Big{]},$
@@
Algorithm. We adapt Stochastic Gradient Ascent and Descent (SGDA) to solve equation 3. SGDA alternates between inner maximization and outer minimization by performing one step of gradient each time. The full algorithm is shown in Appendix A.
Remark: Our method is indeed based on the assumption that the class variable is discrete. Invariance learning methods have primarily been applied to classification tasks where the class variable is typically discrete, which aligns with our requirement. It is worth noting that previous methods often assume that the domains are discrete, which may not be applicable in many applications where the domains are continuous.
@@ 3.1 Theoretical Analysis of Continuous Invariance Learning
In this section, we assume $\bm{t}$ is a one-dimensional scalar to ease the notation for clarity. All results still hold when $\bm{t}$ is a vector. We start by stating some assumptions on the capacity of the function classes.
@@
$\mathcal{H}$ contains $h^{*}$, where $h^{*}(\bm{z}):=\mathbb{E}[\bm{t}|\bm{z}=\Phi(\bm{x})]$.
@@
$\mathcal{G}$ contains $g^{*}$, where $g^{*}(\bm{z},\bm{y}):=\mathbb{E}[\bm{t}|\bm{z}=\Phi(\bm{x}),\bm{y})]$.
@@
Suppose Assumption 2 and 3 hold, $h^{*}$ and $g^{*}$ minimize the following losses given a fixed $\Phi$, we have $h^{*}(\cdot)=\operatorname*{arg\,min}_{h\in\mathcal{H}}\mathbb{E}_{(\bm{x},t)}(h(\Phi(\bm{x}))-\bm{t})^{2},$ and $g^{*}(\cdot)=\operatorname*{arg\,min}_{g\in\mathcal{G}}\mathbb{E}_{(\bm{x},\bm{y},t)}(g(\Phi(\bm{x}),\bm{y})-\bm{t})^{2}.$ where the proof is shown in Appendix G.
@@
Suppose Assumptions 2 and 3 hold. The constraint of equation 3 is satisfied if and only $\mathbb{E}[\bm{t}|\Phi(\bm{x})]=\mathbb{E}[\bm{t}|\Phi(\bm{x}),\bm{y}]$ holds for $\forall\bm{y}\in\mathcal{Y}$.
@@
The advantage of CIL in continuous domain tasks. Consider a 10-class classification task with an infinite number of domains and each domain contains only one sample, i.e., $n\xrightarrow[]{}\infty$, and $n/|\mathcal{T}|=1$, Proposition 1 shows that REx would fail to identify the invariant feature with constant probability. Whereas, Theorem 2 shows that CIL can still effectively extract invariant features. Intuitively, existing methods aim to align $\mathbb{E}^{t}[\bm{y}|\Phi(\bm{x})]$ across different $\bm{t}$ values. However, in continuous environment settings with limited samples per domain, the empirical estimations $\hat{\mathbb{E}}^{t}[\bm{y}|\Phi(\bm{x})]$ become noisy. These estimations deviate significantly from the true $\mathbb{E}^{t}[\bm{y}|\Phi(\bm{x})]$, rendering existing methods ineffective in identifying invariant features. In contrast, CIL proposes to align $\mathbb{E}^{y}[\bm{t}|\Phi(\bm{x})]$, which can be accurately estimated as there are sufficient samples in each class.
We have shown that the definite advantage of CIL based on Proposition 1 and Theorem 2. In the following part, we are going to show the finite sample property of our CIL for completeness. This finite sample analysis is a standard analysis of mini-max formulation based on the results of Lei et al. (2021). We consider the empirical counterpart of the soft regularization version (equation 3) for finite sample performance, i.e.,
$\min_{\bm{w},\Phi,h}\max_{g}\hat{Q}(\bm{w},\Phi,g,h):=\hat{\mathbb{E}}_{\bm{x},\bm{y},\bm{t}}\Big{[}\ell(\bm{w}(\Phi(\bm{x})),\bm{y})+\lambda\big{(}\|h(\Phi(\bm{x}))-\bm{t}\|^{2}_{2}-\|g(\Phi(\bm{x}),\bm{y})-\bm{t}\|_{2}^{2}\big{)}\Big{]}.$
where $\hat{\mathbb{E}}$ is the empirical counterpart of expectation $\mathbb{E}$. Suppose that we solve equation 3.1 with SGDA (Algorithm 1) and obtain $(\hat{w},\hat{\Phi},\hat{g},\hat{h})$, which is a $(\epsilon_{1},\epsilon_{2})$ optimal solution of equation 3.1. Specifically, we have
@@
Suppose we solve equation 3.1 by SGDA as introduced in Algorithm 1 using a training dataset of size $n$ and obtain an $(\epsilon_{1},\epsilon_{2})$ solution $(\hat{w},\hat{\Phi},\hat{g},\hat{h})$. Under the assumptions specified in the appendix, we then have with probability at least $1-\delta$ that